The student answers 3 additional questions when passing the exam. The probability of


answering each question correctly is 0.21. Assuming that all the answers are independent events,


find the probability that at least two correct answers will be given

1
Expert's answer
2022-01-17T18:14:15-0500

The probability that at least two correct answers will be given is:

"P(2,3) = C(2,3)\u00d70.79\u00d70.21^2 + 0.21^3 = 3\u00d70.79\u00d70.21^2 + 0.21^3 = 0.1138" or 11.38%.